integrate the following integral
∫ sin 4x cos 7x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int sin4xsin7xdx
I = \frac{1}{2}\int 2sin7xsin4xdx
I = \frac{1}{2}\int [sin3x+sin10x]dx
I = \frac{1}{2}[\int sin3xdx+\int sin10xdx]
I = \frac{1}{2}[-\frac{cos3x}{3}+\int sin10xdx]
I = \frac{1}{2}[-\frac{cos3x}{3}-\frac{cos10x}{10}]+c